If the ratio of final and initial intensities for 0.1%-solution of L-alanine is 5/8 (length of beaker is 15 mm). determine the ratio of intensities for 0.2% - L-alanine solution in the beaker with length of 23 mm.
1
Expert's answer
2020-10-23T11:54:25-0400

The ratio of final and initial intensities is


"\frac{I_2}{I_1}=\frac{5}8=\frac{\omega k}{d^2}=\frac{0.1k}{15^2},\\\space\\
k=1406.25\text{ mm}^2/\%."

For 0.2% - L-alanine solution in the beaker with length of 23 mm:


"\frac{I_2}{I_1}=\frac{\omega k}{d^2},\\\space\\
\frac{I_2}{I_1}=\frac{0.2 \cdot1406.25}{23^2}=0.53."